OpenStudy (anonymous):

solve by clearing fractions 7/3-5/2x=8+2x ?

sam (.sam.):

(7)/(3)-(5)/(2)*x=8+2x Multiply -(5)/(2) by x to get -(5x)/(2). (7)/(3)-(5x)/(2)=8+2x -(9x)/(2)+(7)/(3)=8 -9x/2=17/3 Multiply each term in the equation by 2 -9x=34/3 Divide each term in the equation by -9 x=-34/27